Jessica has 22 dollars in the bank today. Yesterday, she had 66 dollars in the bank. By what percentage did Jessica's bank account decrease over the past day?
Answer by oscargut(2103)

it decreased 66-22 = 44
44/66 = 2/3 = 0.667
Answer: 66.7%